OpenStudy (anonymous):

0.2x+0.3y=0.4 0.7x+0.4y=0.1 solve for x and y

OpenStudy (texaschic101):

0.2x + 0.3y = 0.4 0.7x + 0.4y = 0.1 if you do not want to mess with decimals, you can multiply both equations by 10 you will then get : 2x + 3y = 4 -->(-4) 7x + 4y = 1-->(3) ------------------ -8x - 12y = -16 (result of multiplying by -4) 21x + 12y = 3 (result of multiplying by 3) -----------------add 13x = -13 -- divide by -13 x = -1 now sub -1 in for x in either equation to find y 2x + 3y = 4 2(-1) + 3y = 4 -2 + 3y = 4 3y = 4 + 2 3y = 6 y = 2 check.. 7x + 4y = 1 7(-1) + 4(2) = 1 -7 + 8 = 1 1 = 1 (correct) x = -1 and y = 2 any questions ?